OpenStudy (anonymous):

7u^2-28m^2 factor the polynomial

OpenStudy (anonymous):

\[7(u ^{2}-4m ^{2})\] \[7(u+2m)(u-2m)\]

OpenStudy (anonymous):

do you understand how to do these type of problems?

OpenStudy (anonymous):

yes im just checking the problems to see if people come up with the same answer as me or something different

OpenStudy (anonymous):

ok, do you have any more?

OpenStudy (anonymous):

i have 7 more

OpenStudy (anonymous):

go ahead, ask away :)

OpenStudy (anonymous):

ok thank you this is the last thats factoring out the rest are solving the equations 16c^2+72cd+81d^2

OpenStudy (anonymous):

this might go faster if you tell me what you got for the answer and i'll check to make sure it's correct

OpenStudy (anonymous):

this one im thinking is prime

OpenStudy (anonymous):

which number do you think is prime?

OpenStudy (anonymous):

im thinking the polynomial is prime like the whole thing in other words it cant be factored down into a product of 2 binomials

OpenStudy (anonymous):

yes. first you should reorder the polynomial so it looks like this:\[72cd+16c ^{2}+81d ^{2}\] then factor \[(4c+9d)(4c+9d)\] \[(4c+9d)^{2}\]

OpenStudy (anonymous):

50b^2+20b+2=0 im thinking the solution to this is 1/5 -1/5

OpenStudy (anonymous):

how did you get the positive 1/5?

OpenStudy (anonymous):

i got the negative one first and the teacher said if the negative works then the positive will work as well and the other way around

OpenStudy (anonymous):

that's only the case when the variaable is squared or has a even number power. try plugging in the positive 1/5 in to the equation. does it come out to equal 0?

OpenStudy (anonymous):

no so it would be -1/5 and what would the other solution be then if its not positive 1/5

OpenStudy (anonymous):

-(1/5) is the only answer. the polynomial should look like this :\[2(25b ^{2}+10b+1)=0\]\[2(5b+1)(5b+1)=0\]\[5b+1=0\]\[b=-{\frac{1}{5}}\]

OpenStudy (anonymous):

sorry my computer is being tarded the next one is g^2+2/3g+1/9=0

OpenStudy (anonymous):

\[g=-{\frac{1}{3}}\]

OpenStudy (anonymous):

okk thanks thats all i needed for now youve been helpful thank you